OpenStudy (anonymous):

how to solve radical equation 5 squared h + 1 = 6 ?

OpenStudy (anonymous):

is it really 5 squared?

OpenStudy (anonymous):

or 5 square root of h+1=6

OpenStudy (anonymous):

its 5 with a radical sign h + 1 = 6

OpenStudy (anonymous):

\[5\sqrt{h+1}=6?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first, divide both sides by 5 \[\sqrt{h+1}=6/5\] square both sides. \[( \sqrt{h+1}=6/5)^{2}\] h+1=36/25 h=36/25-1 h=36/25 - 25/25 h=11/25

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

you are welcome. i hope this helps.